OpenStudy (anonymous):

Help me solve this system of equations? 4x + 3y = -1 3x + y = 3 It's my last question on this test!

OpenStudy (anonymous):

I've solved 10 out of 11 on my own! This is my last one I could use some help on c:

OpenStudy (callisto):

4x + 3y = -1 -(1) 3x + y = 3 => y = 3 - 3x -(2) Sub (2) into (1) 4x + 3 (3 - 3x) = -1 4x + 9 - 9x = -1 -5x = -10 x = ...? Sub x = ... into (2) to find y.

OpenStudy (anonymous):

Or multiply the second equation by (-3) and solve by elimination: 4x+3y=-1 -9x-3y=-9 You get the same thing, -5x=-10 Or graph both equations and find the intersection - so many ways to skin a cat

OpenStudy (anonymous):

friend you can do this sum by using matrix cramers rule.